# atomic orbital API Application Programmer Interface

### v1

@@ Appendix A Löwdin transformation
$\sum_{n=1}^{N}\left\|\psi_{n}-\phi_{n}\right\|^{2}+\sum_{n=1}^{N}\left|b_{n}-a_{n}\right|^{2}.$ (1.1)
@@ A.1. Ground state constrain
on the unknowns $\psi_{n},b_{n}$, which naturally appear in the ground state MCSCF problem. And we want to minimize the distance (1.1) to the given fixed orbitals $\phi_{n}$ and coefficients $a_{n}$. It turns out that the well-known symmetric Löwdin orthonormalisation is the unique solution of this minimization problem [8]. Here we re-derive it making use of the Lagrangian formalism. Firstly, one may easily notice that accounting for the imposed constrain the objective functional can be simplified. In other words, the minimization problem is equivalent to the maximization of
@@
$\psi=S^{-1/2}\phi,\quad b=a/\left\|a\right\|$ (1.2)
@@
It is left to show that the obtained stationary point (1.2) is indeed a maximum. By the Cauchy–Schwarz inequality
@@
for any real vector $b$ with $\left\|b\right\|=1$. Moreover, the equality is achieved if and only if $a,b$ are linearly dependent, which in this case is equivalent to $b=a/\left\|a\right\|$. It shows that the euclidean part of $\mathcal{F}$ is strictly bounded from above by its value at the stationary point (1.2). The corresponding functional inner product part of $\mathcal{F}$ is estimated, firstly, by restricting to the span of the original orbitals $\phi_{1},\ldots,\phi_{N}$. In other words, we suppose that the new orthonormalised functions $\psi_{1},\ldots,\psi_{N}$ can be obtained by a matrix transformation $\psi=\alpha\phi$ of coordinates $\phi_{1},\ldots,\phi_{N}$. Obviously, the matrix $\alpha$ is invertible. From the normalization constrain we deduce
@@
holding true for any vector $x\in\mathbb{R}^{N}$, obviously. This completes the proof of the minimization property of the Löwdin transform (1.2). For an alternative rigorous exposition we refer to [8].
@@ A.2. Excited state constrain
Appendix B Spherical harmonics
$A_{m}(x,y)=\sum_{p=0}^{m}\begin{pmatrix}m\\ p\end{pmatrix}x^{p}y^{m-p}\cos\left((m-p)\frac{\pi}{2}\right)$
$B_{m}(x,y)=\sum_{p=0}^{m}\begin{pmatrix}m\\ p\end{pmatrix}x^{p}y^{m-p}\sin\left((m-p)\frac{\pi}{2}\right)$
@@ Appendix B Spherical harmonics
$\Pi_{l}^{m}(z)=\left(\frac{(l-m)!}{(l+m)!}\right)^{1/2}\sum_{k=0}^{\lfloor(l-m)/2\rfloor}(-1)^{k}2^{-l}\begin{pmatrix}l\\ k\end{pmatrix}\begin{pmatrix}2l-2k\\ l\end{pmatrix}\frac{(l-2k)!}{(l-2k-m)!}r^{2k}z^{l-2k-m}$
Then for $m>0$ we have
$r^{l}\begin{pmatrix}Y_{l,m}\\ Y_{l,-m}\end{pmatrix}=\sqrt{\frac{2l+1}{2\pi}}\Pi_{l}^{m}(z)\begin{pmatrix}A_{m}\\ B_{m}\end{pmatrix}$
and for $m=0$:
$r^{l}Y_{l,0}=\sqrt{\frac{2l+1}{4\pi}}\Pi_{l}^{0}(z)$
Declarations
The CRediT taxonomy of contributor roles [2, 6] is applied. The “Investigation” role also includes the “Methodology”, “Software”, and “Validation” roles. The “Analysis” role also includes the “Formal analysis” and “Visualization” roles. The “Funding acquisition” role also includes the “Resources” role. The contributor roles are visualized in the following authorship attribution matrix, as suggested in Ref. [1].
Table 1. Levels of contribution: major, minor.
ED RVS LF
Conceptualization
Investigation
Data curation
Analysis
Supervision
Writing – original draft
Writing – revisions
Funding acquisition
Project administration

### v2

@@ Appendix A Löwdin transformation
(A.1) $\sum_{n=1}^{N}\left\|\psi_{n}-\phi_{n}\right\|^{2}+\sum_{n=1}^{N}\left|b_{n}-a_{n}\right|^{2}.$
@@ A.1. Ground state constrain
on the unknowns $\psi_{n},b_{n}$, which naturally appear in the ground state MCSCF problem. And we want to minimize the distance (A.1) to the given fixed orbitals $\phi_{n}$ and coefficients $a_{n}$. It turns out that the well-known symmetric Löwdin orthonormalisation is the unique solution of this minimization problem [5]. Here we re-derive it making use of the Lagrangian formalism. Firstly, one may easily notice that accounting for the imposed constrain the objective functional can be simplified. In other words, the minimization problem is equivalent to the maximization of
@@
(A.2) $\psi=S^{-1/2}\phi,\quad b=a/\left\|a\right\|$
@@
It is left to show that the obtained stationary point (A.2) is indeed a maximum. By the Cauchy–Schwarz inequality
@@
for any real vector $b$ with $\left\|b\right\|=1$. Moreover, the equality is achieved if and only if $a,b$ are linearly dependent, which in this case is equivalent to $b=a/\left\|a\right\|$. It shows that the euclidean part of $\mathcal{F}$ is strictly bounded from above by its value at the stationary point (A.2). The corresponding functional inner product part of $\mathcal{F}$ is estimated, firstly, by restricting to the span of the original orbitals $\phi_{1},\ldots,\phi_{N}$. In other words, we suppose that the new orthonormalised functions $\psi_{1},\ldots,\psi_{N}$ can be obtained by a matrix transformation $\psi=\alpha\phi$ of coordinates $\phi_{1},\ldots,\phi_{N}$. Obviously, the matrix $\alpha$ is invertible. From the normalization constrain we deduce
@@
holding true for any vector $x\in\mathbb{R}^{N}$, obviously. This completes the proof of the minimization property of the Löwdin transform (A.2). For an alternative rigorous exposition we refer to [5].
@@ A.2. Excited state constrain
Appendix B Optimality of spin-restricted configuration expansion
Here we prove the existence of closed shell natural expansion (5.1) from first principles. Our proof is valid for both real and complex valued orbitals. In order to clearly distinguish between spatial and spin orbitals in the expressions below, we avoid using Dirac brackets for Slater determinants, as used above, for example, in the expression (5.1). We begin by recalling some basic facts and notations from the electronic structure theory.
Definition 1.
A two-component NN-electron wave function $\Psi$ is said to be non-relativistically closed shell if
(B.1) $\hat{S}^{2}\Psi=0,$
where
(B.2) $\hat{S}^{2}$ $:=\hat{\mathbf{S}}\cdot\hat{\mathbf{S}}$,
(B.3) $\hat{\mathbf{S}}$ $:=\sum_{n=1}^{N}\hat{\mathbf{s}}_{n},$
(B.4) $\hat{\mathbf{s}}_{n}$ $:=1\otimes\dots\otimes\underbrace{\hat{\mathbf{s}}}_{\mathclap{n\textnormal{-th position}}}\otimes\dots\otimes 1,$
(B.8) $\hat{\mathbf{s}}$ $:=\frac{1}{2}\left(\begin{array}[]{c}\sigma_{x}\\ \sigma_{y}\\ \sigma_{z}\end{array}\right)$
and $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ are Pauli matrices.
Remark 1.
$\hat{S}^{2}\Psi=0$ and $\hat{\boldsymbol{S}}\Psi=\boldsymbol{0}$ are equivalent statements because
$\hat{S}^{2}\Psi=0\quad\Longrightarrow\quad\left\langle\Psi\middle|\hat{S}^{2}\Psi\right\rangle=0\quad\Longrightarrow\quad\left\|\hat{S}_{x}\Psi\right\|^{2}+\left\|\hat{S}_{y}\Psi\right\|^{2}+\left\|\hat{S}_{z}\Psi\right\|^{2}=0\quad\Longrightarrow\quad\hat{\mathbf{S}}\Psi=\mathbf{0}.$
Definition 2.
The adjoint of an antilinear operator $\mathcal{V}\overset{\hat{T}}{\longrightarrow}\mathcal{V}$ is the unique antilinear operator $\hat{T}^{\dagger}$ such that
(B.9) $\left\langle\phi_{1}\middle|\hat{T}\phi_{2}\right\rangle=\left\langle\phi_{2}\middle|\hat{T}^{\dagger}\phi_{1}\right\rangle$
for all $\phi_{1},\phi_{2}\in\mathcal{V}$. Antilinearity means that for any scalar $\alpha$ and spin-orbital $\phi\in\mathcal{V}$ it holds that $\hat{T}(\alpha\phi)=\overline{\alpha}\hat{T}\phi$.
Definition 3.
Electronic ladder operators $\bigwedge^{N}\mathcal{V}\overset{\hat{a}_{\phi}}{\longrightarrow}\bigwedge^{N-1}\mathcal{V}$ and $\bigwedge^{N-1}\mathcal{V}\overset{\hat{a}_{\phi}^{\dagger}}{\longrightarrow}\bigwedge^{N}\mathcal{V}$ are defined by
(B.10) $\hat{a}_{\phi}^{\dagger}\Psi={}$ $\phi\wedge\Psi,$
(B.11) $\hat{a}_{\phi}\bigwedge_{n=1}^{N}\psi_{n}={}$ $\sum_{n=1}^{N}(-1)^{n-1}\langle\phi|\psi_{n}\rangle\bigwedge_{n^{\prime}\neq n}\psi_{n^{\prime}},$
where $\mathcal{V}$ denotes the space of spin orbitals. These operators satisfy the anticommutation relations
(B.12) $\{\hat{a}_{\phi_{1}},\hat{a}_{\phi_{2}}\}={}$ $0$
(B.13) $\left\{\hat{a}_{\phi_{1}}^{\dagger},\hat{a}_{\phi_{2}}^{\dagger}\right\}={}$ $0$
(B.14) $\left\{\hat{a}_{\phi_{1}},\hat{a}_{\phi_{2}}^{\dagger}\right\}={}$ $\langle\phi_{1}|\phi_{2}\rangle.$
Theorem 1 (Non-relativistic closed shell two-electron canonical representation).
Let $\Psi$ be a real or complex two-component two-electron wave function of rank $L$, and
$\hat{S}^{2}\Psi=0.$
Then there are real scalars $c_{1},\dots,c_{L}$ and orthonormal orbitals $\psi_{1},\dots,\psi_{L}$ such that
(B.15) $\Psi=\sum_{l=1}^{L}c_{l}\left(\begin{array}[]{c}\psi_{l}\\ 0\end{array}\right)\wedge\left(\begin{array}[]{c}0\\ \psi_{l}\end{array}\right).$
Proof.
The idea of the proof is to associate to $\Psi$ an antilinear operator and to exploit its spectral structure to construct the desired orthonormal orbitals $\psi_{l}$. We begin by defining the following antilinear antihermitian operator
$\mathcal{V}$ $\overset{\hat{T}}{\longrightarrow}\mathcal{V}$
$\psi$ $\mapsto\hat{a}_{\psi}\Psi,$
where $\mathcal{V}$ denotes the inner product space of spin orbitals. For any Hermitian one-electron operator $\hat{O}$ and for any spin orbital $\phi\in\mathcal{V}$, we have
$\left\{\hat{O},\hat{T}\right\}\phi=\hat{a}_{\phi}\hat{O}\Psi.$
Note that $\hat{O}$ can be naturally extended to Fock space $\bigoplus_{N=0}^{\infty}\bigwedge^{N}\mathcal{V}$. This identity can be verified by observing that for all $\phi_{1},\phi_{2}\in\mathcal{V}$ we have
$\left\langle\phi_{1}\middle|\left\{\hat{O},\hat{T}\right\}\phi_{2}\right\rangle={}$ $\left\langle\phi_{1}\middle|\left(\hat{O}\hat{a}_{\phi_{2}}+\hat{a}_{\hat{O}\phi_{2}}\right)\Psi\right\rangle$
$={}$ $\left\langle\left(\hat{a}_{\phi_{2}}^{\dagger}\hat{O}+\hat{a}_{\hat{O}\phi_{2}}^{\dagger}\right)\phi_{1}\middle|\Psi\right\rangle$
$={}$ $\left\langle\left(\phi_{2}\wedge\hat{O}\phi_{1}+\hat{O}\phi_{2}\wedge\phi_{1}\right)\middle|\Psi\right\rangle$
$={}$ $\left\langle\hat{O}\left(\phi_{2}\wedge\phi_{1}\right)\middle|\Psi\right\rangle$
$={}$ $\left\langle\hat{O}\hat{a}_{\phi_{2}}^{\dagger}\phi_{1}\middle|\Psi\right\rangle$
$={}$ $\left\langle\phi_{1}\middle|\hat{a}_{\phi_{2}}\hat{O}\Psi\right\rangle.$
Since $\hat{S}^{2}\Psi=0$, which means $\hat{\boldsymbol{S}}\Psi=\boldsymbol{0}$, and so
$\left\{\hat{\boldsymbol{S}},\hat{T}\right\}=\boldsymbol{0}.$
Since $\Psi$ has rank $L$, it admits a representation
$\Psi=\sum_{l=1}^{L}\phi_{l,1}\wedge\phi_{l,2}$
for some spin orbitals $\{\phi_{l,n}\}_{\begin{subarray}{c}l=1,\dots,L\\ n=1,2\end{subarray}}$. Define the finite-dimensional vector space
$V:=\operatorname{Span}\{\phi_{l,n}\}_{\begin{subarray}{c}l=1,\dots,L\\ n=1,2\end{subarray}}.$
We claim that the spin orbitals $\{\phi_{l,n}\}_{\begin{subarray}{c}l=1,\dots,L\\ n=1,2\end{subarray}}$ are linearly independent, since if they are not, one of the spin orbitals is a linear combination of the others. Without loss of generality we can take that spin orbital to be $\phi_{L,2}$. Then
$\phi_{L,2}=\sum_{l=1}^{L-1}\sum_{n=1}^{2}\alpha_{l,n}\phi_{l,n}+\beta\phi_{L,1}\qquad\text{for some scalars }\{\alpha_{l,n}\}\text{ and }\beta,$
and
$\Psi={}$ $\sum_{l=1}^{L-1}\phi_{l,1}\wedge\phi_{l,2}+\phi_{L,1}\wedge\left(\sum_{l=1}^{L-1}\sum_{n=1}^{2}\alpha_{l,n}\phi_{l,n}+\beta\phi_{L,1}\right)$
$={}$ $\sum_{l=1}^{L-1}\phi_{l,1}\wedge\phi_{l,2}+\alpha_{l,2}\phi_{L,1}\wedge\phi_{l,2}-\alpha_{l,1}\phi_{l,1}\wedge\phi_{L,1}$
$={}$ $\sum_{l=1}^{L-1}(\phi_{l,1}+\alpha_{l,2}\phi_{L,1})\wedge(\phi_{l,2}-\alpha_{l,1}\phi_{L,1}).$
This contradicts the fact that $\operatorname{Rank}\Psi=L$, proving the linear independence. Since
$\hat{T}\psi=\sum_{l=1}^{L}\sum_{n=1}^{2}(-1)^{n-1}\langle\psi|\phi_{l,n}\rangle\bigwedge_{\begin{subarray}{c}n^{\prime}=1\\ n^{\prime}\neq n\end{subarray}}^{2}\phi_{l,n^{\prime}},$
$\operatorname{Im}\hat{T}\subseteq V$, and $\hat{T}$ can be restricted to an operator on $V$. We now show that the restriction
$V\xrightarrow{\hat{T}\big|_{V}}V$
is invertible. It suffices to show that $\operatorname{Ker}\hat{T}\big|_{V}=\{0\}$. Let $\psi\in\operatorname{Ker}\hat{T}\big|_{V}$. Then
$\sum_{l=1}^{L}\langle\psi|\phi_{l,1}\rangle\phi_{l,2}-\langle\psi|\phi_{l,2}\rangle\phi_{l,1}=0.$
Since $\{\phi_{l,n}\}_{\begin{subarray}{c}l=1,\dots,L\\ n=1,2\end{subarray}}$ is linearly independent, $\psi\perp V$ which implies $\psi=0$. Thus $\hat{T}\big|_{V}$ is invertible.
Next, we show that $V$ is an invariant subspace for $\hat{S}_{z}$. Let $\psi\in V$. Then from invertibility of $\hat{T}\big|_{V}$, there is a $\phi\in V$ so that $\psi=\hat{T}\phi$. The anticommutation relation gives us that $V$ is an invariant subspace, since
$\hat{S}_{z}\psi=\hat{S}_{z}\hat{T}\phi=-\hat{T}\hat{S}_{z}\phi\in\operatorname{Im}\hat{T}\subseteq V.$
In particular, the fact that $\hat{S}_{z}$ is diagonalizable in $\mathcal{V}$ with spectrum $\left\{-\frac{1}{2},\frac{1}{2}\right\}$, translates to the restriction $\hat{S}_{z}\big|_{V}$.
The result now follows from Theorem 2, stated and proved below. In order to use Theorem 2, it remains to show that
$\dim E_{-\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)={}$ $\dim E_{\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right),$
$\dim\left(E_{-\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)+E_{\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)\right)={}$ $2\left\lfloor\frac{\dim V}{2}\right\rfloor.$
To show $\dim E_{-\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)=\dim E_{\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)$, we note that
$\hat{T}\left(E_{-\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)\right)\subseteq E_{\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right),$
from the anticommutation relation and use the antilinear isomorphism
$E_{-\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)$ $\longrightarrow E_{\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)$
$\phi$ $\mapsto\hat{T}\phi.$
To show $\dim\left(E_{-\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)+E_{\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)\right)=2\left\lfloor\frac{\dim V}{2}\right\rfloor$, we use the fact that $\hat{S}_{z}\big|_{V}$ is diagonalizable with spectrum $\left\{-\frac{1}{2},\frac{1}{2}\right\}$, which means $V=E_{-\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)\oplus E_{\frac{1}{2}}\left(\hat{S}_{z}\big|_{V}\right)$. Furthermore, $V$ is even-dimensional since $\dim V=2L$. By Theorem 2, there is an orthonormal basis of spin orbitals $\{\psi_{l,n}\}_{\begin{subarray}{c}l=1,\dots,L\\ n=1,2\end{subarray}}$ in $V$, and there are real numbers $c_{1}\dots,c_{L}$ so that for all $l=1,\dots,L$
$\hat{T}\psi_{l,1}={}$ $c_{l}\psi_{l,2},$
$\hat{T}\psi_{l,2}={}$ $-c_{l}\psi_{l,1},$
$\hat{S}_{z}\psi_{l,1}={}$ $\frac{1}{2}\psi_{l,1},$
$\hat{S}_{z}\psi_{l,2}={}$ $-\frac{1}{2}\psi_{l,2}.$
Since the basis is orthonormal, we have that for all $\phi\in V$
$\phi=\sum_{l=1}^{L}\left(\left\langle\psi_{l,1}\middle|\phi\right\rangle\psi_{l,1}+\left\langle\psi_{l,2}\middle|\phi\right\rangle\psi_{l,2}\right).$
Using the special property of the basis, we have that for all $\phi\in V$
$\hat{T}\phi={}$ $\sum_{l=1}^{L}\left(\overline{\left\langle\psi_{l,1}\middle|\phi\right\rangle}\hat{T}\psi_{l,1}+\overline{\left\langle\psi_{l,2}\middle|\phi\right\rangle}\hat{T}\psi_{l,2}\right)$
$={}$ $\sum_{l=1}^{L}c_{l}\left(\left\langle\phi\middle|\psi_{l,1}\right\rangle\psi_{l,2}-\left\langle\phi\middle|\psi_{l,2}\right\rangle\psi_{l,1}\right)$
$={}$ $\hat{a}_{\phi}\sum_{l=1}^{L}c_{l}\psi_{l,1}\wedge\psi_{l,2}.$
We now have two expressions for $\hat{T}\phi$, which gives the equation
$\hat{a}_{\phi}\left(\Psi-\sum_{l=1}^{L}c_{l}\psi_{l,1}\wedge\psi_{l,2}\right)=0\qquad\forall\phi\in V.$
It is straightforward to show that if an NN-electron wave function $\Phi\in\bigwedge^{N}V$ satisfies $\hat{a}_{\phi}\Phi=0$ for all $\phi\in V$ then $\Phi=0$. Consequently,
$\Psi=\sum_{l=1}^{L}c_{l}\psi_{l,1}\wedge\psi_{l,2},$
and we are done up to the proof of Theorem 2. ∎
Proposition 1.
Let $V$ be an inner product space over a field $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $V\overset{\hat{T}}{\longrightarrow}V$ antilinear and
(B.16) $\hat{T}^{\dagger}=-\hat{T}$.
Then the only eigenvalue $\hat{T}$ can have in $\mathbb{K}$ is 0.
Proof.
Let $\lambda\in\mathbb{K}$ and $v\in V\setminus\{0\}$ be an eigenpair. Then
$\hat{T}v={}$ $\lambda v$
$\hat{T}^{2}v={}$ $\overline{\lambda}\hat{T}v$
$\hat{T}^{2}v={}$ $|\lambda|^{2}v$
$\left\langle\hat{T}^{2}v\middle|v\right\rangle={}$ $|\lambda|^{2}\langle v|v\rangle$
$\left\langle\hat{T}^{\dagger}v\middle|\hat{T}v\right\rangle={}$ $|\lambda|^{2}\langle v|v\rangle$
$-\left\langle\hat{T}v\middle|\hat{T}v\right\rangle={}$ $|\lambda|^{2}\langle v|v\rangle.$
Hence $\lambda=0$, since otherwise a positive quantity would equal a negative one. ∎
Theorem 2.
Let $V$ be a finite-dimensional inner product space over the field $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, and let $V\overset{\hat{T}}{\longrightarrow}V$ be an antilinear antihermitian operator, that is
$\hat{T}^{\dagger}=-\hat{T}$,
and let $V\overset{\hat{S}}{\longrightarrow}V$ be a Hermitian $\mathbb{K}$-linear operator with eigenvalues s-s and $s$ so that
$\left\{\hat{T},\hat{S}\right\}$={} $0$
[…]
Then $V$ has an orthonormal basis
$\{u_{1},v_{1},u_{2},v_{2},\dots,u_{N},v_{N}\}$
if it is even-dimensional or
$\{u_{1},v_{1},u_{2},v_{2},\dots,u_{N},v_{N},w\}$
if it is odd-dimensional and there are $\lambda_{1},\dots,\lambda_{N}\in\mathbb{R}$ so that for all $n=1,\dots,N$
$\hat{T}u_{n}={}$ $\lambda_{n}v_{n}$
$\hat{T}v_{n}={}$ $-\lambda_{n}u_{n}$
$\hat{S}u_{n}={}$ $su_{n}$
$\hat{S}v_{n}={}$ $-sv_{n},$
and, in the case where $V$ is odd-dimensional,
$Tw=0.$
Proof.
Use induction on $\left\lfloor\frac{\dim V}{2}\right\rfloor\in\mathbb{N}_{0}$.
$\left\lfloor\frac{\dim V}{2}\right\rfloor=0$:
If $V$ is even-dimensional, $\dim V=0$, and there is nothing to show. If $V$ is odd-dimensional, $\dim V=1$. Then there is a non-zero $w\in V$ and since $Tw\in V$, $Tw=\lambda w$ for some $\lambda\in\mathbb{C}$, which means $Tw=0$ by Proposition 1.
@@ Appendix B Optimality of spin-restricted configuration expansion
Holds for $\left\lfloor\frac{\dim V}{2}\right\rfloor=N$ $\quad\Longrightarrow\quad$ holds for $\left\lfloor\frac{\dim V}{2}\right\rfloor=N+1$:
If $\hat{T}^{2}=0$:
We must have that $\hat{T}=0$ since
$\|Tv\|^{2}=\langle Tv|Tv\rangle=-\left\langle v\middle|T^{2}v\right\rangle,$
which means the theorem is proven by setting all $\lambda_{n}=0$ and picking any orthonormal eigenbasis of $\hat{S}$ since
[…]
which ensures that $N+1$ basis vectors have eigenvalue $s$ and $N+1$ basis vectors have eigenvalue s-s.
If $\hat{T}^{2}\neq 0$:
Since
$\left\{\hat{T},\hat{S}\right\}=0,$
we have
$\left[\hat{T}^{2},\hat{S}\right]=0.$
[…]
$\hat{T}^{2}u_{N+1}=-\mu u_{N+1}.$
Set
$\lambda_{N+1}:={}$ $\sqrt{\mu}$
$v_{N+1}:={}$ $\frac{1}{\lambda_{N+1}}\hat{T}u_{N+1}.$
Then
$\hat{T}u_{N+1}={}$ $\lambda_{N+1}v_{N+1}$
$\hat{T}v_{N+1}={}$ $-\lambda_{N+1}u_{N+1}$
$\|u_{N+1}\|={}$ $\|v_{N+1}\|=1$
$\langle u_{N+1}|v_{N+1}\rangle={}$ $0.$
From the anticommutation relation, we get
$\hat{S}u_{N+1}={}$ $su_{N+1}$
$\hat{S}v_{N+1}={}$ $-sv_{N+1}.$
The subspace $U:=\operatorname{Span}\{u_{N+1},v_{N+1}\}$ is an invariant subspace for $\hat{T}$. The orthogonal complement $U^{\perp}$ is also an invariant subspace for $\hat{T}$ since for all $u\in U$ and $v\in U^{\perp}$
$\left\langle u\middle|\hat{T}v\right\rangle=-\left\langle v\middle|\hat{T}u\right\rangle=0.$
$U^{\perp}$ is an invariant subspace for $\hat{S}$ since $U$ is an invariant subspace for $\hat{S}$ and for all $u\in U$ and $v\in U^{\perp}$
$\left\langle u\middle|\hat{S}v\right\rangle=\left\langle\hat{S}u\middle|v\right\rangle=0.$
To use the induction hypothesis, it remains to show that
$\dim E_{-s}\left(\hat{S}\big|_{U^{\perp}}\right)={}$ $\dim E_{s}\left(\hat{S}\big|_{U^{\perp}}\right)$
$\dim\left(E_{-s}\left(\hat{S}\big|_{U^{\perp}}\right)+E_{s}\left(\hat{S}\big|_{U^{\perp}}\right)\right)={}$ $2N.$
In the case $s\neq 0$, this follows from
[…]
In the case $s=0$, it follows from
$E_{0}\big(\hat{S}\big)=E_{0}\left(\hat{S}\big|_{U^{\perp}}\right)\oplus\operatorname{Span}\{u_{N+1},v_{N+1}\}.$
By the induction hypothesis, $U^{\perp}$ has an orthonormal basis
$\{u_{1},v_{1},u_{2},v_{2},\dots,u_{N},v_{N}\}$
if it is even-dimensional or
$\{u_{1},v_{1},u_{2},v_{2},\dots,u_{N},v_{N},w\}$
if it is odd-dimensional and there are real numbers $\lambda_{1},\dots,\lambda_{n}$ so that for all $n=1,\dots,N$
$Tu_{n}={}$ $\lambda_{n}v_{n}$
$Tv_{n}={}$ $-\lambda_{n}u_{n}$
$\hat{S}u_{n}={}$ $su_{n}$
$\hat{S}v_{n}={}$ $-sv_{n},$
and if $V$ is odd dimensional,
$Tw=0.$
This completes the proof. ∎